OpenStudy (anonymous):

Find the percent of increase from 320 to 380.Round to the nearest 10th if necessary. 18.8% 84.2% 15.8% 0.2% Please I will give you a medal and fan you.please you will get a medal medal.also whenever I SEE YOUR NAME UNDER A ANSWER FROM A QUESTION YOU WILL GET A MEDAL.I'M DESPERATE

OpenStudy (anonymous):

PLEASE PLEASE I''\LL GIVE YOU A MEDAL

hero (hero):

a = 1st Number b = 2nd Number \[\%\Delta = \frac{|a - b|}{a} \times 100\] The percent change equals the absolute value of a minus b divided by a times one hundred.

OpenStudy (anonymous):

OK

hero (hero):

Let me know what you get for the result

hero (hero):

Are you solving it, or do you need a hint?

OpenStudy (anonymous):

A HINT

hero (hero):

\[\%\Delta = \frac{|320 - 380|}{320} \times 100\] You only need to compute the right side of the equal sign.

OpenStudy (anonymous):

Oh i i subtract then i turn my answer into a decimal with percentage

hero (hero):

You have to compute one thing at a time. First compute 320 - 380

OpenStudy (anonymous):

the answer is 18.8%

OpenStudy (anonymous):

well actually i goggled it

hero (hero):

That's pretty sad.

OpenStudy (anonymous):

I know but i have like 5 other subjects to do.

OpenStudy (anonymous):

so I noticed you're still here

hero (hero):

Remember, you're talking to someone who went to High School.

OpenStudy (anonymous):

well my bad.I will be there in 2 years

OpenStudy (anonymous):

do you like my picture i just changed it sorry i took so long to write back.I had to go to Google images to get this.